# Nonlinear bandgap transmission by discrete rogue waves induced in a pendulum chain

### 23 Jan 2024 v1

@@ II.1 Model description
The model under consideration consists of a pendulum chain connected by torsional springs and subjected to a horizontal driving force with frequency $\omega_{d}$ and amplitude $A$, as illustrated in Fig. 1  reproduced from Ref. [37]. Each rigid rod of length $L$ supports the pendulum bob with mass $M$ at its end. The experimental system for the pendulum chain shown in Fig. 1 had been proposed in Ref. [38]. The Lagrangian for a chain of pendula in absence of damping can be written as follows [37]:
$\begin{array}[]{l}\rm\mathcal{L}=\sum\limits_{n=1}^{N}\{\frac{1}{2}I\dot{\theta_{n}}^{2}+\frac{1}{2}\left(Ml+\frac{ml}{2}\right)g\cos(\theta_{n})\\ ~{}~{}~{}~{}~{}~{}~{}+\frac{1}{2}\left(Ml+\frac{ml}{2}\right)\left[2A\omega_{d}\dot{\theta_{n}}\sin(\omega_{d}t_{1})\cos(\theta_{n})\right]\\ ~{}~{}~{}~{}~{}~{}~{}-\frac{1}{4}\beta\left[\left(\theta_{n}-\theta_{n-1}\right)^{2}+\left(\theta_{n}-\theta_{n+1}\right)^{2}\right]\},\end{array}$ (1)
with $I=Ml^{2}+\frac{1}{3}ml^{2}$. The angle $\theta_{n}$ measures the deviation from the vertical for the nth pendulum, $\dot{\theta_{n}}$ is the corresponding angular speed, $g$ denotes the acceleration due to gravity, and $\beta$ is the linear coupling coefficient between pendula due to a torsion spring. The equation of motion for the nth pendulum, derived using the Euler-Lagrange’s equation, is given by [39, 38]:
@@ II.2 Supratransmission threshold
Here, we consider the model without the forcing coefficient, i.e., $f=0$, then Eq. (3) becomes the equation of the Sine-Gordon pendulum. Expanding $\sin(\theta_{n})$ as a Taylor series up to the third order, the equations of motion become
@@
with $a=2+\dfrac{1}{c}\left(1-\omega^{2}\right)$ and $b=\dfrac{{1}}{{8c}}$. This map possesses three fixed points: $x_{0}=0$ and $x_{\pm}=\pm\sqrt{8(1-\omega^{2})}$. The latter two exist only for $\omega<1$. This corresponds to the lower forbidden band and it in agreement with the band for which supratransmission phenomenon has been performed by Geniet and Léon [1]. Fixed points are depicted by a cross in Fig. 2. The stability of a fixed point is obtained by linearizing the map around the fixed point using the procedure described in Ref. [47]. A necessary condition for the existence of manifolds is that the fixed point must be a saddle point. A homoclinic orbit corresponds to an orbit that connects, both in forward and backward time, a saddle fixed point with itself [46]. Within the lower forbidden band ($\omega<1$) only the fixed point $x_{0}=0$ is a saddle point. Figure 2 depicts the progression of the stable (blue line) and unstable (red line) manifolds emanating from the fixed point $x_{0}=0$. The corresponding main homoclinic connection can be clearly observed. The supratransmission threshold corresponds to the value of the turning point of the main homoclinic orbit [15]. Its value can be identified in this work by the dashed green line in Fig. 2.
@@
III Numerical experiment
In this section, numerical studies are carried out on the discrete equation (3) with the ode45 solver of MatLab. This solver uses variable time step to keep the desired relative and absolute tolerance, which are set at $10^{-10}$. The left boundary of the lattice depends on whether or not the lattice is driven. The reflection at the right edge of the lattice will be avoided by choosing a large value of N, but only a few of them will be depicted after full integration of Eq. (3).
III.1 Driven Sine-Gordon pendulum
@@
III.2 Horizontally shaken pendulum
The lattice is subjected to a periodic horizontal displacement of the pivot point with dimensionless frequency $\omega_{1}$ and amplitude $f$. Figure 5 depicts the spatiotemporal evolution of the wave that results from the full integration of Eq. (3) with a shaking dimensionless frequency in the lower forbidden band ($\omega_{1}=0.95$) and the other parameters given by: $f=0.2$ and $c=1$. It is observed, similarly to Ref. [37], that the lattice traps the gap soliton and that, therefore, there is no transmission process. The same phenomenon is observed (not shown here) in the upper forbidden band. The trapping of the gap soliton here is done at different times, and the differences of localization in space are consequence to the staggeringly driving force obtained in Ref. [49]. Another form of localization can be obtained using random fluctuations [50]. By periodically driving the lattice, the gap transmission is observed while the soliton is trapped at a different time when the chain is subjected to a periodic horizontal shaking. Below, the behavior of the chain will be explored when it is subjected to being both simultaneously driven and shaken.
@@ III.2 Horizontally shaken pendulum
III.3 Simultaneously driven and shaken model
Here we have two excitations: periodically driven edges and parametric excitation. The periodically driven edge frequency is taken in the lower forbidden band ($\omega=0.95$). The shaking frequency $\omega_{1}$ will be taken firstly within the linear phonon band ($1\leq\omega\leq\sqrt{1+4c}$) and secondly in the lower forbidden band ($\omega<1$).
Figure 6 shows the spatiotemporal evolution of the chain with a driving frequency ($\omega=0.95$) within the lower forbidden band and driving amplitude ($A=1.22$) below the supratransmission threshold ($A_{thr}=1.255$). The shaking frequency ($\omega_{1}=1.22$) is taken within the linear phonon band. The generation of a train of solitons is observed, although the driving amplitude is below the threshold. Therefore, the presence of the shaking reduces the supratransmission threshold. The same phenomenon is observed (not shown here) when the shaking frequency is taken in the upper forbidden band.
@@ III.3 Simultaneously driven and shaken model
A discrete rogue wave is generated here with a plane wave as the initial condition. To the best of our knowledge, this is the first time that this phenomenon is observed.

### 26 Jan 2024 v2

@@ II.1 Model description
The model under consideration consists of a pendulum chain connected by torsional springs and subjected to a horizontal driving force with frequency $\omega_{d}$ and amplitude $A$, as illustrated in Fig. 1  reproduced from Ref. [37]. Each rigid rod of length $l$ and mass $m$ supports the pendulum bob with mass $M$ at its end. The experimental system for the pendulum chain shown in Fig. 1 had been proposed in Ref. [38]. The Lagrangian for a chain of pendula in absence of damping can be written as follows [37]:
$\begin{array}[]{l}\mathcal{L}=\sum\limits_{n=1}^{N}\{\dfrac{1}{2}I\dot{\theta_{n}}^{2}+\dfrac{1}{2}\left(Ml+\dfrac{ml}{2}\right)g\cos(\theta_{n})\\ ~{}~{}~{}~{}~{}~{}~{}+\dfrac{1}{2}\left(Ml+\dfrac{ml}{2}\right)\left[2A\omega_{d}\dot{\theta_{n}}\sin(\omega_{d}t_{1})\cos(\theta_{n})\right]\\ ~{}~{}~{}~{}~{}~{}~{}-\dfrac{1}{4}\beta\left[\left(\theta_{n}-\theta_{n-1}\right)^{2}+\left(\theta_{n}-\theta_{n+1}\right)^{2}\right]\},\end{array}$ (1)
with $I=Ml^{2}+\frac{1}{3}ml^{2}$ being the system moment of inertia. The angle $\theta_{n}$ measures the deviation from the vertical for the nth pendulum, $\dot{\theta_{n}}$ is the corresponding angular speed, $g$ denotes the acceleration due to gravity, and $\beta$ is the linear coupling coefficient between pendula due to a torsion spring. The equation of motion for the nth pendulum, derived using the Euler-Lagrange’s equation, is given by [39, 38]:
@@ II.2 Supratransmission threshold
Here, we consider the model without the forcing coefficient, i.e., $f=0$, then Eq. (3) becomes the discrete Sine-Gordon equation. Expanding $\sin(\theta_{n})$ as a Taylor series up to the third order, the equations of motion become
@@
with $a=2+\dfrac{1}{c}\left(1-\omega^{2}\right)$ and $b=\dfrac{{1}}{{8c}}$. This map possesses three fixed points: $x_{0}=0$ and $x_{\pm}=\pm\sqrt{8(1-\omega^{2})}$. The latter two exist only for $\omega<1$. This frequency band corresponds to the lower forbidden band, which is in agreement with the band for which the supratransmission phenomenon has been observed by Geniet and Léon [1]. Fixed points are depicted by a cross in Fig. 2. The stability of a fixed point is obtained by linearizing the map around the fixed point using the procedure described in Ref. [47]. A necessary condition for the existence of manifolds is that the fixed point must be a saddle point. A homoclinic orbit corresponds to an orbit that connects, both in forward and backward time, a saddle fixed point with itself [46]. Within the lower forbidden band ($\omega<1$) only the fixed point $x_{0}=0$ is a saddle point. Figure 2 depicts the progression of the stable (blue line) and unstable (red line) manifolds emanating from the fixed point $x_{0}=0$. Stable and unstable manifolds emanating from the fixed point $x_{0}=0$ intersect and form a homoclinic orbit, which can be identified by the loop in Fig. 2. The first-order connections of the loop correspond to the main homoclinic orbit as can be clearly observed in Fig. 2. The supratransmission threshold corresponds to the value of the turning point of the main homoclinic orbit [15]. Its value can be identified in this work by the dashed green line in Fig. 2.
@@
III Numerical experiments
In this section, numerical studies are carried out on the discrete equation (3) with the ode45 solver of MatLab. This solver uses variable time step to keep the desired relative and absolute tolerance, which are set at $10^{-10}$. The left boundary of the lattice depends on whether or not the lattice is driven. The reflection at the right edge of the lattice will be avoided by choosing a large value of N=101 and appropriate time for the full integration of Eq. (3).
III.1 Driven pendulum chain without shaking
@@
III.2 Horizontally shaken pendulum chain
The lattice is subjected to a periodic horizontal displacement of the pivot point with dimensionless frequency $\omega_{1}$ and amplitude $f$. Figure 5 depicts the spatiotemporal evolution of the wave that results from the full integration of Eq. (3) with a shaking dimensionless frequency in the lower forbidden band ($\omega_{1}=0.95$) and the other parameters given by: $f=0.2$ and $c=1$. It is observed, similarly to Ref. [37], that the lattice traps the gap soliton as and that, therefore, there is no transmission process. The same phenomenon is observed (not shown here) in the upper forbidden band.
The trapping of the gap soliton here is an analogy with the light being trapped into several neighboring waveguides obtained in Ref. [14], but it is done at different times. The trapping here is different from the localization in space, which is a consequence of the staggeringly driving force obtained in Ref. [49]. Another form of localization can be obtained using random fluctuations [50]. By periodically driving the lattice, the gap transmission is observed while the soliton is trapped at a different time when the chain is subjected to a periodic horizontal shaking. Below, the behavior of the chain will be explored when it is subjected to being both simultaneously driven and shaken.
@@ III.2 Horizontally shaken pendulum chain
III.3 Simultaneously driven and shaken pendulum chain
Here we have two excitations: periodically driven edge and parametric excitation. The periodically driven edge frequency is taken in the lower forbidden band ($\omega=0.95$). The shaking frequency $\omega_{1}$ will be taken firstly within the linear phonon band ($1\leq\omega\leq\sqrt{1+4c}$) and secondly in the lower forbidden band ($\omega<1$).
Figure 6 shows the spatiotemporal evolution of the chain with a driving frequency ($\omega=0.95$) within the lower forbidden band and with driving amplitude ($A=1.22$) below the supratransmission threshold ($A_{thr}=1.255$). The shaking frequency ($\omega_{1}=1.22$) is taken within the linear phonon band. The generation of a train of solitons is observed, although the driving amplitude is below the threshold. Therefore, the presence of the shaking reduces the supratransmission threshold. The same phenomenon is observed (not shown here) when the shaking frequency is taken in the upper forbidden band.
@@ III.3 Simultaneously driven and shaken pendulum chain
A discrete rogue wave is generated here with a plane wave as the initial condition. To the best of our knowledge, this is the first time that this phenomenon is observed. It is worth pointing out that presently it is not possible to discard that a rogue wave can also be generated in the just driven, not shaken case, although it has not yet been observed.
III.4 Spectral analysis
IV CONCLUSIONS